# Classification of unimodal isolated complete intersection singularities in positive characteristic

### v2

@@ Abstract.
In this paper, we classify unimodal isolated complete intersection singularities over fields of positive characteristic under contact equivalence. This work extends the classification of simple isolated complete intersection singularities in positive characteristic, which was established by T.H. Pham, G. Pfister, and G. M. Greuel. To achieve this classification, we extend the complete transversal method to fields of positive characteristic, a generalization that is also applicable to numerous other classification problems.
@@
algebraic geometry, singularity theory
@@
Zuo is supported by NSFC Grant 12271280 and BNFS Grant 1252009
@@ 1. Introduction
Classification is one of the oldest topic in singularity theory. The modality of singularities for real and complex hypersurfaces are first introduced by V. I. Arnold in [AVGZ12]. He also finished the classification of hypersurface singularities with small modality over $\mathbb{C}$ in [Arn76]. G.-M. Greuel and H.D. Nguyen generalized the notion of modality to the algebraic setting in [GN16], so that one can define modality over any algebraically closed field of arbitrary characteristic. They also classified simple (i.e. modality 0) hypersurface singularities in positive characteristic field under right equivalence.
@@
The classification of isolated complete intersection singularities (ICIS) under contact equivalence was studied in the 1980s. Assume $(X,0)=(f^{-1}(0),0)$ be a complete intersection germ with an isolated singularity defined by $f:K^{n}\rightarrow K^{p}$. We call such germs $I_{n,p}$. M. Giusti has shown that only $I_{1,1}$, $I_{2,2}$ and $I_{3,2}$ can be simple in [Giu83]. He then classified all simple ICIS over $K=\mathbb{C}$. The classification of unimodal (i.e. modality 1) germs from plane to plane ($I_{2,2}$ case) was completed by A. Dimca and C.G. Gibson. C.T.C. Wall then classified the unimodal germs of $I_{n,p}$ with $n>p$. Recently, T.H. Pham, G. Pfister, and G.-M. Greuel generalized the modality of hypersurface singularities to ICIS, and they classified simple zero-dimensional ICIS ($I_{2,2}$ case) over any algebraically closed field of arbitrary characteristic. In this paper, we continue their work. We use a different method to classify unimodal zero-dimensional ICIS of $I_{2,2}$ over any algebraically closed field of arbitrary characteristic. In Corollary 2.24, we have shown that for all zero-dimensional ICIS, only $I_{2,2}$ and $I_{3,3}$ can be unimodal. Unfortunately, the classification of $I_{3,3}$ seems very complicated and needs more new tools.
@@
In the previous work [BAM20], the finite determinacy theorem works well in zero characteristic fields. T.H. Pham and G.-M. Greuel generalized it to positive characteristic in [PG19]. Although the given boundary is sharp, it is still insufficient to deal with complicated (i.e. higher order) problems. For example, let $h=(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4}+y^{5})$ be the $5$-jet of an ICIS $f$ over an algebraically closed field $F$. If $char(F)=0$, then $h$ is $5$-determined. Therefore $f$ is contact equivalent to $(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4}+y^{5})$, and then to $(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4})$ after a transformation. Otherwise, we can only know that $h$ is $7$-determined in positive characteristic, which gives few idea about the form of $f$.
@@
To solve the problem, we develop new tools. We generalized the complete transversal method introduced in [BKdP99] into positive characteristic field, which is useful in many classification problem, see Corollary 2.17. It can be used for semi-quasi homogeneous singularities over field of arbitrary characteristic. For the above example, $f$ is a semi-quasi homogeneous singularity with initial term $(x^{3}+xy^{3},x^{2}y+y^{4})$. Using our method, we can show that $f$ is contact equivalent to $(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4})$.
@@
The main result is Theorem 8.4. Surprisingly, the classification result in positive characteristic turns out to be similar to the zero characteristic case except for some special characteristic.
@@ 2. Basic settings
We first recall some basic concepts of ICIS. Then we introduce the finite determinacy in positive characteristic in [PG19]. After that, we generalize the complete transversal method to positive characteristic. The estimation of modality is also given in this section.
@@ 2.1. Basic concepts
(2) We call $f=(f_{1},\dots,f_{m})$ an isolated complete intersection singularity (ICIS) if $I=\langle f_{1},\dots,f_{m}\rangle$ defines a complete intersection and there exists $k\in\mathbb{N}$ such that $\mathfrak{m}^{k}\subset I+I_{m}(J(f))$, where $J(f)=(\frac{\partial f_{i}}{\partial x_{j}})_{ij}$ is the $m\times n$ Jacobian matrix and $I_{m}(J(f))$ is the ideals generated by all $m\times m$ minors of $J(f)$. Denote
@@
In this article, we mainly discuss $I_{2,2},$ which denotes the zero-dimensional isolated complete intersection singularity in plane.
@@
and the action of $\mathcal{\mathcal{K}}$ acting on $R^{m}$ is defined as
@@
where $\phi(\mathbf{x})=(\phi(\mathbf{x_{1}}),\dots,\phi(\mathbf{x_{n}}))$.
Let $f$ and $g$ define two isolated complete intersections with same codimension $n-m$. $f$ is called contact equivalent to $g$, denoted by $f\sim g$, if $g\in\mathcal{K}f$, i.e. there exists $U\in GL(m,R)$ and $\phi\in Aut(R)$ such that $g=U\cdot\phi(f)$. By [GP19], $f\sim g$ if and only if $\langle f_{1},\dots,f_{m}\rangle\sim\langle g_{1},\dots,g_{m}\rangle$, i.e. there exists $\phi\in Aut(R)$ such that $\phi(\langle f_{1},\dots,f_{m}\rangle)=\langle g_{1},\dots,g_{m}\rangle$.
@@ 2.2. Tangent image and finite determinacy
(1) The $k$jet space of $R^{m}$ is defined as $J_{k}=R^{m}/\mathfrak{m}^{k+1}R^{m}$. For $f\in R^{m}$, the $k$-jet of $f$ is the image in $J_{k}$, denoted by $j_{k}(f)$. Let $\pi:J_{l}\rightarrow J_{k}$ be the natural projection and denote the kernel as $P_{k,l}.$ If $f\in J_{k}$ is a $k$-jet, we denote the submanifold $J_{l}(f)=f+P_{k,l}$.
@@
Let $J_{k}=R^{m}/\mathfrak{m}^{k+1}R^{m}$ denote the $k$jet space of $R^{m}$. For $f\in R^{m}$, the $k$-jet of $f$ is the image in $J_{k}$, denoted by $j_{k}(f)$. Let $\mathcal{K}_{k}=\{(j_{k}(U),j_{k}(\phi))\mid U\in GL(m,R),\ \phi\in Aut(R)\}$ be the $k$-jet algebraic group and the algebraic action of $\mathcal{K}_{k}$ on affine space $J_{k}$ is defined as
@@
[…]
(2.1) $\widetilde{T}_{f}(\mathcal{K}_{k}f)=T_{f}(\mathcal{K}_{k}f)$
[…]
The tangent image are also computed in [PG19] as following.
@@
The finite determinacy are strongly related to tangent image.
@@
Let $f$ be an isolated complete intersection singularity, then $f$ is $22\tau(f)-ord(f)+$ 2-determined.
@@ 2.3. Complete transversal
Let $f$ be a $k$-jet in $J_{k}$ and weighted homogeneous of degree 00 w.r.t. $(a;d)$. Moreover, assume
(2.2) $\mathrm{max}(d_{i})<(k+1)\mathrm{min}(a_{j})\mathrm{\ or\ }\mathrm{min}(d_{i})>(k+1)\mathrm{max}(a_{j})$
For $l>k$, let $P_{k,l},\ J_{l}(f)$ be the subset of $J_{l}$ defined in Definition 2.4.(1). We have the following useful theorem from [BAM20].
@@
We can mainly refer to the proof of [BAM20] Proposition 1.3. Although the original proof is based on $\mathbb{C}$, nothing has changed in positive characteristic. ∎
@@
Condition 2.9 is necessary. In [BGM10] Remark 1 there is a counterexample without 2.9.
@@ 2.4. Complete transversal and homogeneous filtrations
Next we introduction a filtration of contact group $\mathcal{K}$ compatible with the weighted filtration. For details one may see [BKdP99] Section 2.3.
@@
For the survey of the standard Mather group $\mathcal{K},\mathcal{R},\mathcal{C}$, one can refer [Mat68].
@@
(i) $F^{r}\mathcal{K}$ respects the filtration $\{F_{a,d}^{r}R^{m}\}$, i.e. for $(U,\phi)\in F^{r}\mathcal{K},\ f\in F^{s}R^{m},\ U\cdot\phi(f)\in F^{s}R^{m}$.
@@
(iii) The Lie algebra algebra action satisfies the following: for any $f-g\in F_{a,d}^{t}R^{m}$ with $f,g\in F_{a,d}^{0}R^{m}$ and $l\in Lie(F_{a,d}^{r}R^{m})$, we have $l\cdot f-l\cdot g\in F_{a,d}^{r+t}R^{m}.$
@@
(2.3) $\widetilde{T}_{f}(F_{a,d}^{r}\mathcal{K}\cdot f)=F_{a\cup d,d}^{r}(\langle f_{1},\dots,f_{m}\rangle\cdot R^{m})+\sum_{j}F_{a,a_{j}}^{r}(\mathfrak{m})\cdot\frac{\partial f}{\partial x_{j}}.$
@@
(2.4) $f(x+z)$ $=f(x)+\sum_{i=1}^{n}\xi_{i}\frac{\partial f}{\partial x_{i}}+$
@@
Set $W=F_{a,d}^{k+1}R^{m}\backslash F_{a,d}^{k+2}R^{m}$ be the subspace of $F_{a,d}^{0}R^{m}$. We only need to show the following hold for any $f\in F_{a,d}^{0}R^{m}$ and $w\in W$:
@@
(ii) $f+\{Lie(F^{1}\mathcal{K})\cdot x\cap W\}\subset F^{1}\mathcal{K}\cdot x\cap\{x+W\}+F_{a,d}^{k+2}R^{m}.$
Then the rest part is same as the proof of [BKdP99] Theorem 2.25.
(i) follows from Proposition 2.12.(iii). For (ii), let $w=(w_{1},\dots,w_{m})\in\{Lie(F^{1}\mathcal{K})\cdot x\cap W\}$. Then
@@
by Lemma 2.14, where $h_{i}$ are the higher order terms defined in 2.4. We have $h_{i}\in F_{a,d}^{k+2}R^{m}$ for $i=1,2,3$ since $\xi_{i}$ appear twice in each terms and $v_{a}(\xi_{i})\geq v_{a}(x_{i})+k+1$. We also have $Uw\in F_{a,d}^{k+2}R^{m}$ since for each $j$, $v_{a}(\lambda_{ji}w_{i})=v_{a}(\lambda_{ji})+v_{a}(w_{i})\leq d_{j}-d_{i}+k+1+d_{i}+k+1\geq d_{j}+k+2$.
Set $\widetilde{w}=h_{1}+h_{2}+h_{3}-Uw$. By above discussion, $\widetilde{w}\in F_{a,d}^{k+2}R^{m}.$ Without loss of generality, we can assume $v_{a}(f_{1})\leq\dots\leq v_{a}(f_{m}).$ Then $U(0)$ is an upper triangular matrix whose principal diagonal element is zero. Hence $Id-U$ must be invertible. We have $f+w\sim(Id-U)\cdot\phi(f+w)=f+\widetilde{w}.$ Denote $g=(Id-U,\phi)\in\mathcal{K}$, we have $f+w=g^{-1}(f)+g^{-1}(\widetilde{w})$.
To finish (ii), it remains to show $g\in F^{1}\mathcal{K}$. Since $\xi=(\xi_{1},\dots,\xi_{n})\in F_{a,a}^{1}R^{n}$ we have $\phi\in I_{n}+F_{a,a}^{1}R^{n}=F^{1}\mathcal{R}$ by Definition 2.10. Similarly $Id-U\in I_{m}+F_{a\cup d,a\cup d}^{1}R^{m}\subset F^{1}\mathcal{C}.$ Now we get $g=(Id-U,\phi)\in F^{1}\mathcal{K}$. ∎
@@
In [BKdP99], Mather’s lemma implies (i) $\Rightarrow$ (ii). But the proof of Mather’s lemma depends on analysis in complex field. We complete the proof in this special case without Mather’s lemma. The proof refers Marc Giusti’s proof of [Giu83] Proposition 1.
Use induction, we can show
@@
Suppose $f$ is weighted homogeneous of weight $r$ w.r.t. $(a;d)$. Take $s>r$. Let $T$ be a subspace of $F_{a,d}^{r+1}R^{m}$ such that $F_{a,d}^{r+1}R^{m}\subset T+\widetilde{T}_{f}(F_{a,d}^{1}\mathcal{K}\cdot f)+F_{a,d}^{s+1}R^{m}$, then any $g$ with $g-f\in F_{a,d}^{r+1}R^{m}$ is $F^{1}\mathcal{K}$-equivalent to $f+t+\phi$ where $t\in T$ and $\phi\in F_{a,d}^{s+1}R^{m}$.
@@
Suppose $f$ is an ICIS of weight $r$ w.r.t. $(a,d)$. Let $T$ be a subspace of $F_{a,d}^{r+1}R^{m}$ such that

### v3

@@ Abstract.
In this paper we classify the unimodal isolated complete intersection singularities in arbitrary characteristic under contact equivalence. The classification over $\mathbb{C}$ has already done by A. Dimca and C.G. Gibson. We continue and generalize their work. To complete the classification, we generalized the complete transversal method into positive characteristic field, which is also useful in many other classification problem.
@@
Unimodal singularities, complete intersection, positive characteristic
@@
Zuo is supported by NSFC Grant 12271280 and BJNSF Grant 1252009. Yau is supported by the Tsinghua University Education Foundation.
@@ 1. Introduction
Classification is one of the oldest topic in singularity theory. The modality of singularities for real and complex hypersurfaces are first introduced by V.I. Arnold in [AVGZ12]. He also finished the classification of hypersurface singularities with small modality over $\mathbb{C}$ in [Arn76]. G.M. Greuel and H.D. Nguyen generalized the notion of modality to the algebraic setting in [GN16], so that one can define modality over any algebraically closed field of arbitrary characteristic. They also classified simple (i.e. modality 0) hypersurface singularities in positive characteristic field under right equivalence.
@@
The classification of isolated complete intersection singularities (ICIS) under contact equivalence was studied in the 1980s. Assume $(X,0)=(f^{-1}(0),0)$ to be a complete intersection germ with an isolated singularity defined by $f:F^{n}\rightarrow F^{p}$. We call such germs $I_{n,p}$. M. Giusti has shown that only $I_{1,1}$, $I_{2,2}$ and $I_{3,2}$ can be simple in [Giu83]. He then classified all simple ICIS over $F=\mathbb{C}$. The classification of unimodal (i.e. modality 1) germs from plane to plane ($I_{2,2}$ case) was completed by A. Dimca and C.G. Gibson. C.T.C. Wall then classified the unimodal germs of $I_{n,p}$ with $n>p$. Recently, T.H. Pham, G. Pfister, and G.M. Greuel generalized the modality of hypersurface singularities to ICIS, and they classified zero-dimensional ICIS ($I_{2,2}$ case) over any algebraically closed field of arbitrary characteristic. In this paper, we continue their work. We use a different method to classify unimodal zero-dimensional ICIS of $I_{2,2}$ over any algebraically closed field of arbitrary characteristic. In Corollary 2.24, we have shown that for all zero-dimensional ICIS, only $I_{2,2}$ and $I_{3,3}$ can be unimodal. Unfortunately, the classification of $I_{3,3}$ seems very complicated and needs more new tools.
@@
In the previous work [DG83], the finite determinacy theorem works well in zero characteristic fields. T.H. Pham and G.M. Greuel generalized it to positive characteristic in [PG19]. Although the given boundary is sharp, it is still insufficient to deal with complicated (i.e. higher order) problems. For example, let $h=(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4}+y^{5})$ be the $5$-jet of an ICIS $f$ over an algebraically closed field $F$. If $char(F)=0$, then $h$ is $5$-determined. Therefore $f$ is contact equivalent to $(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4}+y^{5})$, and then to $(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4})$ after a transformation. Otherwise, we can only know that $h$ is $7$-determined in positive characteristic, which gives few idea about the form of $f$.
@@
To solve the problem, we develop new tools. We generalized the complete transversal method introduced in [BKdP99] into positive characteristic field, which is useful in many classification problem, see Corollary 2.17. It can be used for semi-quasi homogeneous singularities over field of arbitrary characteristic. For the above example, $f$ is a semi-quasi homogeneous singularity with initial term $(x^{3}+xy^{3},x^{2}y+y^{4})$. Using our method we can show that $f$ is contact equivalent to $(x^{3}+xy^{3}+y^{5},x^{2}y+y^{4})$.
@@
The main result is Theorem 8.4. Surprisingly, the classification result in positive characteristic turns out to be similar as the zero characteristic case excepting for some specical characteristic.
@@ 2. Basic settings
We first talk about some basic concepts of ICIS. Then we introduce the finite determinacy in positive characteristic in [PG19]. After that, we generalize the complete transversal method into positive characteristic. The estimation of modality is also given in this section.
@@ 2.1. Basic concepts
(2) We call $f=(f_{1},\dots,f_{m})$ an isolated complete intersection singularity (ICIS) if $I=\langle f_{1},\dots,f_{m}\rangle$ defines a complete intersection and there exists $k\in\mathbb{N}$ such that $\mathfrak{m}^{k}\subset I+I_{m}(J(f))$, where $J(f)=(\frac{\partial f_{i}}{\partial x_{j}})_{ij}$ is the $m\times n$ Jacobian matrix and $I_{m}(J(f))$ is the ideal generated by all $m\times m$ minors of $J(f)$. Denote
@@
In this article, we focus on $I_{2,2},$ which denotes the zero-dimensional isolated complete intersection singularity in the plane.
@@
and the action of $\mathcal{\mathcal{K}}$ on $R^{m}$ is defined as
@@
where $\phi(\mathbf{x})=(\phi(x_{1}),\dots,\phi(x_{n}))$.
Let $f$ and $g$ define two isolated complete intersections of the same codimension nmn-m. $f$ is called contact equivalent to $g$, denoted by $f\sim g$, if $g\in\mathcal{K}f$, i.e. there exists $U\in GL(m,R)$ and $\phi\in Aut(R)$ such that $g=U\cdot\phi(f)$.
@@ 2.2. Tangent image and finite determinacy
(1) The $k$ -jet space of $R^{m}$ is defined as $J_{k}=R^{m}/\mathfrak{m}^{k+1}R^{m}$. For $f\in R^{m}$, the $k$-jet of $f$ is the image in $J_{k}$, denoted by $j_{k}(f)$. Let $\pi:J_{l}\rightarrow J_{k}$ be the natural projection and denote the kernel as $P_{k,l}.$ If $f\in J_{k}$ is a $k$-jet, we denote the submanifold $J_{l}(f)=f+P_{k,l}$.
@@
Let $J_{k}=R^{m}/\mathfrak{m}^{k+1}R^{m}$ denote the $k$ -jet space of $R^{m}$. Let $\mathcal{K}_{k}=\{(j_{k}(U),j_{k}(\phi))\mid U\in GL(m,R),\ \phi\in Aut(R)\}$ be the $k$-jet algebraic group and the algebraic action of $\mathcal{K}_{k}$ on affine space $J_{k}$ is defined as
@@
The tangent space $T_{e}(\mathcal{K}_{k})$ of algebraic group $\mathcal{K}_{k}$ has a natural Lie algebra structure (see [WR05] chapter 4). The orbit map $\pi:\mathcal{K}_{k}\rightarrow\mathcal{K}_{k}\cdot f$ induces the tangent map $d\pi:Lie(\mathcal{K}_{k})\rightarrow T_{f}(\mathcal{K}_{k}f)$. We denote the image of $d\pi$ as $\widetilde{T}_{f}(\mathcal{K}_{k}f)$, which coinsides with $Lie(\mathcal{K}_{k})\cdot f$. When charF=0F=0, $d\pi$ is surjective and
$\widetilde{T}_{f}(\mathcal{K}_{k}f)=T_{f}(\mathcal{K}_{k}f)$ (2.1)
(and therefore $\widetilde{T}_{f}(\mathcal{K}f)=T_{f}(\mathcal{K}f)$). But when charF>0F>0, (2.1) may not hold. For details one can see [PG19], section 2.
The tangent image is computed in [PG19] as follows.
@@
The finite determinacy is strongly related to tangent image.
@@
Let $f$ be an isolated complete intersection singularity, then $f$ is ($2\tau(f)-ord(f)+2$)-determined.
@@ 2.3. Complete transversal
Let $f$ be a $k$-jet in $J_{k}$ and weighted homogeneous of degree 0 with respect to $(a;d)$. Moreover, assume
$\mathrm{max}(d_{i})<(k+1)\mathrm{min}(a_{j})\mathrm{\ or\ }\mathrm{min}(d_{i})>(k+1)\mathrm{max}(a_{j}).$ (2.2)
For $l>k$, let $P_{k,l},\ J_{l}(f)$ be the subsets of $J_{l}$ defined in Definition 2.4.(1). We have the following useful theorem from [DG83].
@@
We primarily refer to the proof of Proposition 1.3 in [DG83]. Although the original proof is based on $\mathbb{C}$, nothing needs to be changed in positive characteristic. ∎
@@
Condition (2.2) is necessary. In [BGM12] Remark 1 there is a counterexample without (2.2): let $f=y^{2}+xy^{3}\in F[[x,y]]$ with charF=2F=2. One can check that $f$ does not satisfies condition $2.$ Computation shows $\widetilde{T}_{f}(\mathcal{K}_{l}f)=\mathfrak{m}^{4}/\mathfrak{m}^{l}$, which implies the complete transversal $C=0$ for any $l>4$ and then for any $g\in f+\mathfrak{m}^{5}$, $g\sim f$. But in fact we have $f+x^{5}$ is not contact equivalent to $f$.
@@ 2.4. Complete transversal and homogeneous filtrations
Next we introduce a filtration of contact group $\mathcal{K}$ compatible with the weighted filtration. For details one may see [BKdP99] Section 2.3.
@@
For the survey of the standard Mather group $\mathcal{K},\mathcal{R},\mathcal{C}$, one can refer to [Mat68].
@@
(i) $F^{r}\mathcal{K}$ respects the filtration $\{F_{a,d}^{r}R^{m}\}$, i.e. for every $r,s\geq 0$, $(U,\phi)\in F^{r}\mathcal{K},\ f\in F^{s}R^{m},\ U\cdot\phi(f)\in F^{s}R^{m}$, where $F^{s}R^{m}=F^{s}_{a,d}R^{m}$ is the submodule of $R^{m}$ generated by the monomials of degree equal or greater than $s$.
@@
(iii) The Lie algebra action satisfies the following: for any $f-g\in F_{a,d}^{t}R^{m}$ with $f,g\in F_{a,d}^{0}R^{m}$ and $l\in Lie(F_{a,d}^{r}R^{m})$, we have $l\cdot f-l\cdot g\in F_{a,d}^{r+t}R^{m}.$
@@
$\widetilde{T}_{f}(F_{a,d}^{r}\mathcal{K}\cdot f)=F_{a\cup d,d}^{r}(\langle f_{1},\dots,f_{m}\rangle\cdot R^{m})+\sum_{j}F_{a,a_{j}}^{r}(\mathfrak{m})\cdot\frac{\partial f}{\partial x_{j}},$ (2.3)
@@
${f(x+\xi)}$ $=f(x)+\sum_{i=1}^{n}\xi_{i}\frac{\partial f}{\partial x_{i}}+$ (2.4)
@@
Set $W=F_{a,d}^{k+1}R^{m}\backslash F_{a,d}^{k+2}R^{m}$ be the subspace of $F_{a,d}^{0}R^{m}$.
Claim: the following hold for any $f\in F_{a,d}^{0}R^{m}$ and $w\in W$:
@@
(ii) $f+\{Lie(F^{1}\mathcal{K})\cdot f\cap W\}+F_{a,d}^{k+2}R^{m}\subset F^{1}\mathcal{K}\cdot f\cap\{f+W\}+F_{a,d}^{k+2}R^{m}.$
Proof of claim:
(i) follows from Proposition 2.12 (iii). For (ii), let $w=(w_{1},\dots,w_{m})\in\{Lie(F^{1}\mathcal{K})\cdot f\cap W\}$. Then
@@
by Lemma 2.14, where $h_{i}$ are the higher order terms defined in (2.4). We have $h_{i}\in F_{a,d}^{k+2}R^{m}$ for $i=1,2,3$ since $\xi_{i}$ appear twice in each terms and $v_{a}(\xi_{i})\geq v_{a}(x_{i})+k+1$. We also have $Uw\in F_{a,d}^{k+2}R^{m}$ since for each $j$, $v_{a}(\lambda_{ji}w_{i})=v_{a}(\lambda_{ji})+v_{a}(w_{i}){\color[rgb]{0,0,0}\geq}d_{j}-d_{i}+k+1+d_{i}+k+1\geq d_{j}+k+2$.
[…]
To finish (ii), it remains to show $g\in F^{1}\mathcal{K}$. Since $\xi=(\xi_{1},\dots,\xi_{n})\in F_{a,a}^{1}R^{n}$ we have $\phi\in I_{n}+F_{a,a}^{1}R^{n}=F^{1}\mathcal{R}$ by Definition 2.10. Similarly $Id-U\in I_{m}+F_{a\cup d,a\cup d}^{1}R^{m}\subset F^{1}\mathcal{C}.$ Now we get $g=(Id-U,\phi)\in F^{1}\mathcal{K}$. The claim is proved.
We note that by claim (i) and (ii),
$\bigcup_{t\in T}F^{1}\mathcal{K}\cdot(f+t+F_{a,d}^{k+2}R^{m})$ $\supset\bigcup_{t\in T}\{f+t+Lie(F^{1}\mathcal{K})\cdot(f+t)\cap W+F_{a,d}^{k+2}R^{m}\}$ (2.5)
$=\bigcup_{t\in T}\{f+t+Lie(F^{1}\mathcal{K})\cdot f\cap W\}+F_{a,d}^{k+2}R^{m}$
$=f+T+Lie(F^{1}\mathcal{K})\cdot f\cap W+F_{a,d}^{k+2}R^{m}$
$=f+(T+Lie(F^{1}\mathcal{K})\cdot f)\cap W+F_{a,d}^{k+2}R^{m}$
$=f+W+F_{a,d}^{k+2}R^{m}.$
That is, for any $g=f+w+\bar{g}$ with $w\in W$ and $\bar{g}\in F_{a,d}^{k+2}R^{m}$, $g$ is contact equivalent to $f+t+\bar{f}$ with $t\in T$ and $\bar{f}\in F_{a,d}^{k+2}R^{m}$. ∎
@@
In [BKdP99], Mather’s lemma implies (i) $\Rightarrow$ (ii). However, the proof of Mather’s lemma relies on analysis in the complex field. We complete the proof in this special case without Mather’s lemma. The proof refers M. Giusti’s proof of [Giu83] Proposition 1.
Using induction, we can show
@@
Suppose that $f$ is weighted homogeneous of weight $r$ with respect to $(a;d)$. Take $s>r$. Let $T$ be a subspace of $F_{a,d}^{r+1}R^{m}$ such that
$F_{a,d}^{r+1}R^{m}\subset T+\widetilde{T}_{f}(F_{a,d}^{1}\mathcal{K}\cdot f)+F_{a,d}^{s+1}R^{m},$
then any $g$ with $g-f\in F_{a,d}^{r+1}R^{m}$ is $F^{1}\mathcal{K}$-equivalent to $f+t+\phi$ where $t\in T$ and $\phi\in F_{a,d}^{s+1}R^{m}$.
@@
Suppose that $f$ is an ICIS of weight $r$ with respect to $(a,d)$. Let $T$ be a subspace of $F_{a,d}^{r+1}R^{m}$ such that